8. A photographer in a helicopter ascending vertically at a constant rate of 1.2 m/s accidentally drops m above the ground. (a) How long will it take the camera to reach the ground? a camera out the window when the helicopter is 23.0 -21.2 9. (b) What will its speed be just before it hits the ground? m/s
Answer #1

9)

V^2 = u^2 + 2as

V^2 = 1.2^2 + 2 x 9.8 x 23

V = 21.27 m/s

8)

V = u + at

-21.27 = 1.2 - 9.8 t

Time, t = 2.05 sec

Comment in case any doubt please rate my answer....